# PINN-MEP: Continuous Neural Representations for Minimum-Energy Path Discovery in Molecular Systems

### v5

@@ Abstract
Characterizing conformational transitions in physical systems remains a fundamental challenge, as traditional sampling methods struggle with the high-dimensional nature of molecular systems and high energy barriers between stable states. These rare events often represent the most biologically significant processes, like ion channel opening that controls cellular signaling, yet may require months of continuous simulation to observe. We present a method that reformulates transition path generation as a continuous optimization problem solved through physics-informed neural networks (PINNs) inspired by string methods for minimum-energy path (MEP) generation. By representing paths as implicit neural functions and leveraging automatic differentiation with differentiable molecular force fields, our method efficiently discovers physically realistic transition pathways without expensive sampling. We demonstrate our approach on two proteins, including an explicitly hydrated BPTI system with more than 8,300 atoms. Our method identifies a representative minimum-energy path for the conformational change captured in a millisecond-scale MD simulation (Shaw et al., 2010), while requiring only force field evaluations compared to approximately 412 billion in the original study. Though our method yields a single pathway rather than an ensemble of transition trajectories, this represents a six orders of magnitude reduction in computational cost, generating a physically meaningful transition in less than 15 minutes on a standard GPU rather than weeks on specialized hardware.
@@ 1 Introduction
Characterizing molecular conformational changes, or transitions, is essential for understanding biological processes, from enzyme catalysis to protein folding. Although molecular dynamics (MD) simulations provide unprecedented atomic-level detail of these processes (Dror et al., 2012), capturing rare transition events remains computationally challenging due to the inherent timescale gap between molecular motion and conformational changes. This difficulty arises from transition rates between stable states being exponentially suppressed by the height of the energy barriers that separate them, as described by transition state theory (Eyring, 1935). As a result, direct simulation methods spend the vast majority of computational effort sampling already known stable states, making the observation of transitions between them prohibitively expensive.
@@
The search for efficient methods to identify transition pathways between stable states has a rich history in computational physics. Early theoretical work established key principles through the Freidlin-Wentzell (Freidlin & Wentzell, 1998) and Onsager-Machlup (Onsager & Machlup, 1953) functionals, which connect the most probable trajectories, including transitions, to minimizers of specific action functionals. Building on these foundations, the minimum-energy path (MEP) between stable molecular states represents the most probable transition pathway in the limit of high friction. This path reveals the mechanism of conformational change, identifies key transition states and intermediate configurations, and can thus be highly informative about the system’s dynamics. Moreover, the MEP can serve as a scaffold for subsequent sampling of the full transition path ensemble, making it a powerful tool for understanding molecular dynamics. Methods such as the string method family or chain-of-states approaches (E et al., 2002; 2005; Ren et al., 2005; Maragliano et al., 2006; E et al., 2007; Petersen et al., 2024; Dellago et al., 1998) approximate these paths as splines or as a discrete series of molecular configurations connected by artificial spring forces to ensure continuity. Despite their strong theoretical foundation, these approaches face fundamental challenges stemming from their optimization in Cartesian molecular coordinates. The discrete chain of conformations requires periodic reparameterization to maintain even spacing, leading to a complex alternation between energy minimization and geometric constraints. Additional limitations include the difficulty of generating good initial transition guesses and limited parallelization, which restricts the scale of practical applications. While this formulation has proven useful for small-scale systems, it has not yet benefited from modern optimization approaches that could address these challenges.
Recent advances in machine learning methodology have created opportunities for novel approaches to this challenge. Physics-informed neural networks (PINNs) have increased our ability to solve complex physical problems by embedding differential equations directly into neural architectures (Raissi et al., 2019). Additionally, the development of differentiable molecular force fields (Wang et al., 2023) has enabled seamless integration of physical force field models used in MD simulations with modern machine learning frameworks, allowing efficient computation of derivatives beyond forces, which are needed for neural path optimization.
@@
Numerous approaches have emerged over the past decades to tackle the problem of sampling transitions. Traditional simulation-based sampling methods (Jung et al., 2023; Lazzeri et al., 2023; Bolhuis et al., 2002; Dellago et al., 1998; Laio & Parrinello, 2002; Torrie & Valleau, 1977) rely on expensive molecular dynamics simulations, often requiring initial MD-harvested transition trajectories or prior knowledge of the transition dynamics to begin with. More recent machine learning approaches have explored several directions: variational approaches using Doob’s h-transform (Du et al., 2024; Lee et al., 2025; Das & Limmer, 2019; Singh & Limmer, 2023), stochastic control formulations (Yan et al., 2022; Holdijk et al., 2023), and even attempts to sample transition trajectories directly using diffusion models (Petersen et al., 2023; Han et al., 2024; Jing et al., 2024).
@@
Notably, Du et al. (2024) uses a similar neural representation of transition paths but with a different objective, targeting the complete transition path ensemble by additionally parameterizing path width and optimizing based on a framework rooted in Doob’s h-transform (Doob, 1957) and Wasserstein Lagrangian flows (Neklyudov et al., 2024). Most closely related to our method, Ramakrishnan et al. (2025) employ a continuous neural parameterization of MEPs inspired by Nudged Elastic Band (NEB) methods, a class of methods related to string methods, incorporating specialized components for tangential forces, path velocity, and transition state energetics to optimize the neural path. Another promising concurrent approach by Raja et al. (2025) creatively leverages pretrained diffusion models as a sampler of initial transition conformations and as a force field to generate plausible transition paths using the discretized Onsager-Machlup action directly as a loss function. While each approach offers valuable insights into neural network-based path discovery, our method focuses on a streamlined optimization scheme that efficiently scales to much larger biomolecular systems with explicit solvent, essential for accurately modeling such systems, while maintaining fast training times.
Drawing inspiration from neural implicit representations in computer vision (Mildenhall et al., 2020; Sitzmann et al., 2020), we reformulate the traditional minimum-energy-path finding problem as a continuous optimization framework. Our approach transforms the Onsager-Machlup functional, which naturally models the probability of paths in stochastic systems, into a remarkably simple and intuitive loss function that requires only energy evaluations along the neural network parametrized path. In contrast to existing methods that use complex composite losses involving energies, forces, and path gradients, our streamlined formulation eliminates the need for these additional computations. This simplification, while faster, also sidesteps known issues in PINN optimization if the loss contains differential operators (Rathore et al., 2024). This PINN framework efficiently captures physically meaningful transition mechanisms for complex biomolecular systems, completing optimization in minutes on standard hardware.
@@ 2.1 Derivation of the Energy-Based Loss Function and Neural Implementation
To derive our loss function, which when minimized yields a continuous curve representing the MEP connecting fixed endpoints and , we build upon alternative formulations of the Onsager-Machlup action functional as explored by Vanden-Eijnden & Heymann (2008) and Olender & Elber (1997) in the context of string methods. However, we further simplify the approach after reaching their geometric formulation. A complete derivation with all mathematical details is provided in Appendix A.1. We begin with a molecular system evolving under overdamped Langevin dynamics, a model for molecular motion in the high-friction regime characteristic of biomolecular systems:
@@
Thus, maximizing the probability of a transition path (subject to fixed endpoints and ) is equivalent to minimizing this action functional. Notably, there are different derivations of the action that yield an additional higher-order derivative term related to trajectory entropy, whose necessity is conditional on the task (Adib, 2008). Since we are primarily interested in MEPs, we neglect this term. However, the resultant MEPs still empirically align well with the free energy landscape. To further simplify, we expand the integrand of the action functional:
@@
Since and are fixed in our formulation, this term becomes a constant that doesn’t affect the optimization. We can therefore focus on minimizing the remaining terms. Applying the Cauchy-Schwarz inequality to the remaining terms:
@@
While the referenced works (Vanden-Eijnden & Heymann, 2008; Olender & Elber, 1997) derive and use this geometric action, we further simplify it. This geometric integral represents a natural endpoint for theoretical development for traditional chain-of-states and string methods. In these classical approaches, forces directly update molecular conformations to minimize , making minimization along a discretized string using forces conceptually straightforward.
@@
where and .
For a path segment, applying Taylor’s theorem, we can express the energy difference as:
Where is the angle between the gradient and displacement vectors. For MEPs, the displacement aligns with or against the gradient direction () giving:
This gives us a surrogate objective with the same minimum:
@@
where are uniformly or randomly sampled points in . By minimizing , we drive the path representation to discover low-energy regions connecting stable states, approximating the MEP predicted by the Onsager-Machlup framework.
2.1.1 Neural Representation of MEPs
We propose representing the path using a neural parameterization, similar to the mean parameterization of Du et al. (Du et al., 2024). While our approaches share similar representations, they differ in objectives and computational efficiency (see Appendix B.1 for theoretical connections and performance comparisons, respectively). Specifically, we construct a continuous function mapping a progress coordinate to molecular configurations , where is the number of atoms, as illustrated in Figure 1 (left):
Here is a neural network, and are basis functions that satisfy the endpoint constraints , , , and , ensuring that and regardless of the neural network’s output. While simple linear functions and can be used, other basis functions with appropriate boundary conditions are also effective. The blending function controls the neural network’s contribution along the path, typically designed to allow maximum influence in the transition region while smoothly vanishing at endpoints (e.g., ).
2.1.2 Path Optimization
We minimize using the Adam optimizer (Kingma & Ba, 2017):
Unlike traditional chain-of-states methods, explicit spring-like continuity terms are unnecessary; the inherent continuity of , combined with smooth blending functions, ensures path continuity.
@@ 2.2 Established Approaches to MEP Finding
Classical approaches to MEP discovery operate directly in Cartesian atomic coordinates, predating differentiable molecular force fields (Wang et al., 2023) that enable our neural network representation. This direct optimization faces significant challenges as steric clashes create extreme energy barriers that are difficult to overcome with local optimization. The string method and chain-of-states approaches discussed below represent broad families with numerous variations, extensively cataloged in reviews such as by E & Vanden-Eijnden (2010). These established techniques illustrate both foundational concepts and limitations that our approach addresses.
2.2.1 String Method
The string method (E et al., 2002; 2005; Ren et al., 2005; Maragliano et al., 2006; E et al., 2007) determines the MEPs by evolving a curve in the configuration space according to the physical forces while maintaining proper parameterization. Given the same endpoints and used in our formulation, the method seeks a curve (string) parameterized using splines mapping to the systems Cartesian coordinates that connects these states. The MEP satisfies:
@@ 2.2.1 String Method
where is the component of the force perpendicular to the unit tangent vector along the string. This orthogonality condition has an intuitive interpretation: at each point along the MEP, the forces acting on the system are entirely parallel to the path itself, with no perpendicular components trying to push the system away from the path. This represents a ridge or valley in the energy landscape—the system naturally follows the path without being forced sideways.
The simplest dynamics for evolving a curve toward this MEP is given by:
Here is a Lagrange multiplier term enforcing the parameterization constraint, typically equal arc-length parameterization where (constant) for all . In practice, the string is discretized into images and evolved through a two-step procedure that requires explicit reparameterization steps:
1. Evolution step:
2. Reparameterization step to enforce equal spacing:
This two-step procedure converges to a discretized MEP satisfying each image’s local orthogonality condition .
2.2.2 Chain-of-States Methods
An alternative perspective views MEP finding as optimization of a chain-of-states (Pratt, 1986; Elber & Karplus, 1987; Ulitsky & Elber, 1990; Jónsson et al., 1998; Henkelman et al., 2000; Gillilan & Lilien, 2004; Sheppard et al., 2008). Given a sequence with and , the objective function is defined as:
where and is a spring constant. The spring term enforces the continuity of the path while allowing flexibility in the discretization. However, this formulation introduces a well-known limitation: the spring forces tend to cause "corner cutting," where the path takes shortcuts through high-energy regions rather than following the low-energy valleys representing physically meaningful transitions. Although variants of chain-of-states methods exist that avoid this limitation, they come at an increased computational cost.

### v6

# Follow the MEP: Scalable Neural Representations for Minimum-Energy Path Discovery in Molecular Systems

@@ Abstract
Characterizing conformational transitions in physical systems remains a fundamental challenge, as traditional sampling methods struggle with the high-dimensional nature of molecular systems and high-energy barriers between stable states. These rare events often represent the most biologically significant processes, yet may require months of continuous simulation to observe. One way to understand the function and mechanics of such systems is through the minimum energy path (MEP), which represents the most probable transition pathway between stable states in the high-friction, low-temperature limit. We present a method that reformulates MEP discovery as a fast and scalable neural optimization problem. By representing paths as implicit neural representations and training with differentiable molecular force fields, our method discovers transition pathways without expensive sampling. Our approach scales to large biomolecular systems through a simple loss function derived from the path’s likelihood via the Onsager-Machlup action and a scalable new architecture, AdaPath. We demonstrate this approach on two proteins, including an explicitly hydrated BPTI system with more than 3,500 atoms. Our method identifies a MEP that captures the same conformational change observed in a millisecond-scale molecular dynamics (MD) simulation in just minutes on a standard GPU, rather than weeks on a specialized cluster.
@@ 1 Introduction
When studying the function of large biological systems, molecular dynamics simulations effectively serve as a computational microscope, providing atomic-level detail and control that are impossible with experimental methods alone (Dror et al., 2012). Their true value often lies in not exhaustive sampling of the system’s Boltzmann distribution—which is intractable for such systems—but in generating hypotheses from single observed transitions that can then be corroborated through targeted in vitro or in vivo experiments. However, capturing even these individual transition events remains computationally challenging due to the inherent timescale gap between molecular motion and conformational changes. This difficulty arises from transition rates between stable states being exponentially suppressed by the height of the energy barriers that separate them, as described by transition state theory (Eyring, 1935). As a result, direct simulation methods spend the vast majority of computational effort sampling already known stable states, making the observation of transitions between them prohibitively expensive, often requiring months of continuous simulation to observe a single biologically significant event.
@@
The search for efficient methods to identify transition pathways between stable states has a rich history in computational physics. Early theoretical work established key principles through the Freidlin-Wentzell (Freidlin & Wentzell, 1998) and Onsager-Machlup (Onsager & Machlup, 1953) functionals, which connect the most probable trajectories, including transitions, to minimizers of specific action functionals. Building on these foundations, the minimum-energy path (MEP) between stable molecular states represents the most probable transition pathway in the limit of high friction and low temperature. This path reveals the mechanism of conformational change, identifies key transition states and intermediate configurations, and can thus be highly informative about a system’s dynamics. Moreover, the MEP can serve as a scaffold for subsequent sampling of the whole transition path ensemble, making it a powerful tool for understanding molecular dynamics. Methods such as the string method family or chain-of-states approaches (E et al., 2002; 2005; Ren et al., 2005; Maragliano et al., 2006; E et al., 2007; Petersen et al., 2024; Dellago et al., 1998) approximate these paths as a discrete series of molecular configurations connected either by artificial spring forces or held together by frequent reparametrization during optimization to ensure continuity. Despite their strong theoretical foundation, these approaches face fundamental challenges stemming from their optimization in Cartesian molecular coordinates (Ovchinnikov et al., 2011). The discrete chain of conformations requires either slow periodic reparameterization or a spring force that introduces a hyperparameter to maintain even spacing (Lindgren et al., 2019). Additional limitations include the difficulty of generating good initial transition guesses (Ovchinnikov et al., 2011) and limited parallelization, which restricts the scale of practical applications. While this formulation has proven useful for small-scale systems, it has not yet benefited from modern optimization approaches that could address these challenges.
Recent advances in machine learning have opened the door to strategies addressing this problem. Physics-informed neural networks (PINNs) have shown that neural networks can be trained directly on physical laws expressed as differential equations (Raissi et al., 2019). In addition, the emergence of differentiable molecular dynamics force fields (Wang et al., 2023) has extended this principle to molecular simulations. Together, these developments provide a foundation for deep learning or other optimization approaches that require efficient computation of derivatives beyond forces.
@@
Numerous approaches have emerged over the past decades to tackle the problem of sampling transitions. Traditional simulation-based sampling methods (Jung et al., 2023; Lazzeri et al., 2023; Bolhuis et al., 2002; Dellago et al., 1998; Laio & Parrinello, 2002; Torrie & Valleau, 1977) rely on expensive molecular dynamics simulations, often requiring initial MD-harvested transition trajectories or prior knowledge of the transition dynamics. More recent machine learning approaches have explored several directions: variational approaches using Doob’s h-transform (Du et al., 2024; Lee et al., 2025; Das & Limmer, 2019; Singh & Limmer, 2023), stochastic control formulations (Yan et al., 2022; Holdijk et al., 2023), and even attempts to sample transition trajectories directly using diffusion models (Petersen et al., 2023; Han et al., 2024; Jing et al., 2024).
@@
Notably, Du et al. (2024) also uses a neural representation of transition paths but with a different objective, targeting the complete transition path ensemble by additionally parameterizing path width and optimizing based on a framework rooted in Doob’s h-transform (Doob, 1957) and Wasserstein Lagrangian flows (Neklyudov et al., 2024). Most closely related to our method, Ramakrishnan et al. (2025) employs a continuous neural parameterization of MEPs inspired by Nudged Elastic Band (NEB) methods, incorporating specialized components for tangential forces, path velocity, and transition state energetics. Another concurrent approach by Raja et al. (2025) leverages pretrained diffusion models as both samplers and force fields. However, these neural approaches face two key limitations: they employ standard MLP architectures that scale poorly with system size or large pretrained models, and they rely on complex multi-term loss functions involving higher-order differential terms known to hamper optimization (Rathore et al., 2024).
Drawing inspiration from neural implicit representations in computer vision (Mildenhall et al., 2020; Sitzmann et al., 2020), we aim to address both limitations. We derive a simple, single-term energy-based loss directly from the Onsager-Machlup functional and introduce AdaPath, a novel architecture that employs parameter sharing across atoms. Unlike existing approaches that enforce continuity through complex loss terms, our neural representation achieves smooth paths inherently through architectural design and optimizer choice. This combination enables efficient scaling to large biomolecular systems while maintaining fast, stable optimization. We further perform extensive multi-system benchmarks of our method against classical and deep learning approaches, and find that our approach achieves far lower path and transition state energies.
@@ 2.1 Derivation of the Energy-Based Loss Function and Neural Implementation
To derive our loss function, which, when minimized, yields a continuous curve representing the MEP connecting fixed endpoints and , we build upon alternative formulations of the Onsager-Machlup action functional as explored by Vanden-Eijnden & Heymann (2008) and Olender & Elber (1997) in the context of string methods. However, we further simplify the approach after reaching their geometric formulation. A complete derivation with all mathematical details is provided in Appendix A.1. To introduce the Onsager-Machlup action intuitively, we begin with a molecular system evolving under overdamped Langevin dynamics, a model for molecular motion in the high-friction regime characteristic of biomolecular systems:
@@
Thus, maximizing the probability of a transition path (subject to fixed endpoints and ) is equivalent to minimizing this action functional. Notably, there are different derivations of the action that yield an additional higher-order derivative term related to trajectory entropy, whose necessity is conditional on the task (Adib, 2008). Since we are primarily interested in MEPs, we neglect this term. However, the resultant MEPs still empirically align well with the free energy landscape. Notably, this might not be the case for strongly entropy-driven systems such as intrinsically disordered proteins, among others (Granata et al., 2015; Caro et al., 2017; Skriver et al., 2023). To further simplify, we expand the integrand of the action functional:
@@
Since and are fixed in our formulation, this term becomes a constant that does not affect the optimization. We can therefore focus on minimizing the remaining terms. Applying the Cauchy-Schwarz inequality to the remaining terms:
@@
While the referenced works (Vanden-Eijnden & Heymann, 2008; Olender & Elber, 1997) derive and use this geometric action, we further simplify it. This geometric integral represents a natural endpoint for theoretical development for traditional chain-of-states and string methods. In these classical approaches, forces directly update molecular conformations to minimize , making minimization along a discretized string conceptually straightforward.
@@
where and . For a path segment, applying Taylor’s theorem, we can express the energy difference as:
Where is the angle between the gradient and displacement vectors. For MEPs, the displacement aligns with or against the gradient direction (). Critically, this approximation requires to be sufficiently small for the higher-order term to be negligible. This yields:
This gives us a surrogate objective with the same minimum, provided the Taylor series approximation holds:
@@
where are uniformly sampled points in . While this energy-based loss successfully identifies MEPs when the underlying assumptions hold, it admits a pathological minimum when they break down. Specifically, if is large, the validity of the Taylor series approximation breaks down. In that case, the loss can be minimized by paths that remain in one stable state and then rapidly ”teleport” to the other, avoiding the high-energy transition region. To prevent this pathological solution, we employ specific architectural choices and optimizer selection to control that we detail in the following section.
2.1.1 AdaPath: A Scalable Neural Architecture for MEPs
Neural approaches to modeling transitions map directly to , requiring hidden layers that scale with system dimensionality, leading to quadratic parameter scaling that limits applicability to large systems.
AdaPath 1 addresses this through parameter sharing across atoms. Each atom has a learnable embedding , processed through shared transformer-like MLP blocks with adaptive layer norm (adaLN) conditioning (Peebles & Xie, 2023). A conditioning network processes sinusoidal embeddings of to generate scale , shift , and gate parameters that modulate each MLP block , and in addition to a final scale and shift before the final projection layer that projects to the atoms’ Cartesian coordinates .
Neural implicit representations (NIR) effectively function as smooth lookup tables with a fixed range of inputs rather than complex mappings as in conventional deep learning tasks, allowing neural networks to remain small while embeddings drive the parameter count. In AdaPath, this enables conditioning parameters, which are computed once per value, to be heavily parameterized, while keeping the main branch operating on each atom embedding lightweight.
AdaPath maintains endpoint constraints through:
where and provide linear interpolation ensuring and , and controls network contribution, vanishing at endpoints.
The residual MLP design maintains path continuity and enables gradient flow. While our energy-based loss does not explicitly enforce continuity, residual connections preserve smooth transitions by allowing direct information flow from embeddings to output. However, standard optimizers can cause excessive path stretching around transition regions, see 6. We therefore use the Muon optimizer for small systems (Keller Jordan et al., 2024), as it maintains lower Lipschitz constants (Large et al., 2024), ensuring adequate path presence in transition regions for accurate energy estimates. For large systems, Adam (Kingma & Ba, 2017) suffices; we posit that this might be due to neural networks’ increasing inability to model discontinuities as dimensionality increases (Petersen & Voigtlaender, 2018). We further scale the energy function and use gradient clipping, as outlined in H, to avoid high energy spikes caused by steric clashes between atoms.
@@ 2.2 Established Approaches to MEP Finding
All MEP discovery methods must solve a fundamental optimization challenge: finding paths that simultaneously minimize energy while maintaining smooth continuity between endpoint configurations. This dual objective creates an inherent tension; energy minimization alone may produce discontinuous jumps between low-energy regions, while continuity enforcement alone may force paths through high-energy barriers. Different methodological families resolve this through distinct mathematical formulations, each with characteristic advantages and limitations that our neural approach addresses. The approaches discussed below represent broad families with numerous variations, extensively cataloged in reviews such as E & Vanden-Eijnden (2010).
2.2.1 Explicit Continuity Penalties
The chain-of-states family (Pratt, 1986; Elber & Karplus, 1987; Ulitsky & Elber, 1990) handles the energy-continuity trade-off through composite objective functions that explicitly balance both terms. Given a sequence with and , the optimization targets:
Where the first term enforces energy minimization and the spring term constant penalizes discontinuities. However, this formulation suffers from ”corner cutting”, where the path takes shortcuts through high-energy regions when spring forces dominate.
The Nudged Elastic Band (NEB) method (Jónsson et al., 1998; Henkelman et al., 2000) refines this approach by projecting spring forces onto the path normal direction, yielding:
Where represents forces perpendicular to the path and the spring term acts only parallel to it. This decoupling reduces corner cutting but requires careful force projection and still faces scaling challenges in high-dimensional systems, often leading to jagged paths (Zhang et al., 2016).
2.2.2 Continuity via Geometric Constraints
The string method (E et al., 2002; 2005; Ren et al., 2005; Maragliano et al., 2006) takes an alternative approach, separating energy minimization from continuity enforcement through alternating optimization. The method seeks a curve satisfying the orthogonality condition:
@@ 2.2.2 Continuity via Geometric Constraints
where is the unit tangent vector along the string. This orthogonality condition has an intuitive interpretation: at each point along the MEP, the forces acting on the system are entirely parallel to the path itself, with no perpendicular components trying to push the system away from the path. The simplest dynamic for evolving toward this MEP is the following:
where the Lagrange multiplier enforces parameterization constraints. However, in practice, implementations alternate between: (1) evolving images along the string using energy gradients , and (2) reparameterization to maintain equal spacing:
This effectively enforces the same constraint as the Lagrange multiplier (E & Vanden-Eijnden, 2010).
2.2.3 Computational Limitations
These classical approaches share fundamental optimization limitations stemming from their discrete optimization in Cartesian molecular coordinates (Ovchinnikov et al., 2011). Chain-of-states methods require careful tuning of the spring constant and face corner cutting, while string methods demand complex reparameterization procedures (Lindgren et al., 2019). Both struggle with initial transition guesses (Ovchinnikov et al., 2011), local minima trapping, and limited parallelization that restricts applications to smaller systems, as each image is separately parameterized. Our neural representation addresses these challenges by encoding continuity architecturally, simplifying optimization to a single energy-based objective while leveraging optimization in neural parameter space to reduce local minima susceptibility.